One type of gasohol is 10% alcohol and another type is 35% alcohol. How much of type must be mixed to obtain 400 gallons of a 25% alcohol mixture?
Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

.10X+.35(400-X)=400*.25
.10X+140-.35X=100
-.25X=-40
X=-40/-.25
X=160 GASOHOL @ 10%
400-160=240 GASOHOL OF 35%
PROOF
.10*160+.35*240=100
16+84=100
100=100

Answer by stanbon(75887)

One type of gasohol is 10% alcohol and another type is 35% alcohol. How much of type must be mixed to obtain 400 gallons of a 25% alcohol mixture?
---------------
Let amount of 10% solution be "x"; amount of alcohol is 0.10x
Amount of 35% solution is "400-x"; amount of alcohol is 0.35(400-x)=140-0.35x
EQUATION:
alcohol + alcohol = 0.25(400)
0.10x+140-0.35x = 100
-0.25x=-40
x=160 gallons (amount of 10% solution)
400-x=240 gallons (amount of 35% solution)
---------
Cheers,
Stan H.